Here is my question.
Jack and John mix two kinds of feed for cats. They wish to make 24 pounds of feed worth $0.45 per pound by mixing one kind worth $0.34 per pound with another worth $0.57 per pound. How many pounds of the cheaper kind should they use in the mix.
Thank for you help.

Answer by Paul(988)

34x+57(24-x)=45(24)
34x+1368-57x=1080
-23x=-288
x=12.5
24-12.5=11.5
Hence, they should mix 12.5 pounds of mix worth 0.34.